Consider a Poisson distribution with a mean of 4. What is the probability that X < 3?

.238
0
.433
None of these answers are correct
Answer #1

Solution :

Given that ,

mean = = 4

Using Poisson probability formula,

P(X = x) = (e- * x ) / x!

P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)

= (e-4 * 40) / 0! + (e-4 * 41) / 1! + (e-4 * 42) / 2! + (e-4 * 43) / 3!

= 0.238

Probability = 0.238